# Adaptive action supervision in reinforcement learning from real-world multi-agent demonstrations

### 15 Dec 2023 v3

@@ 5.1 Performance on Chase-and-escape
First, we verified our method using a chase-and-escape task, in which the predators and prey interacted in a two-dimensional world with continuous space and discrete time. The numbers of predators and prey were 2 and 1, respectively. We first describe the common setting between the source (demonstration) and target RL tasks. The environment was constructed by modifying an environment called the predator-prey in Multi-Agent Particle Environment (MAPE) [lowe2017multi, tsutsui2022emergence, tsutsui2022collaborative]. Following [tsutsui2022collaborative], the play area size was constrained to the range of -1 to 1 on the $x$ and $y$ axes, all agent (predator/prey) disk diameters were set to 0.1, landmarks (obstacles) were eliminated, and predator-to-predator contact was ignored for simplicity. The predators were rewarded for capturing the prey ($+1$), namely contacting the disks, and punished for moving out of the area ($-1$), and the prey was punished for being captured by the predator or for moving out of the area ($-1$).
Fig. 2a shows an example of the chase-and-escape task. The time step was 0.1 s and the time limit in each episode was set to 30 s. The initial position of each episode was randomly selected from a range of -0.5 to 0.5 on the $x$ and $y$ axes. If the predator captured the prey within the time limit, the predator was successful; otherwise, the prey was successful. If one side (predators/prey) moved out of the area, the other side (prey/predators) was deemed successful. There are 13 actions including acceleration in 12 directions at every 30 degrees in relative coordinate system and doing nothing. For the relative mobility of predators and prey in both environments, to examine the effect of domain adaptation, we set the same mobility of the prey in the source and target RL environments, but for the predators, we set 140% and 107% of the prey mobility in the source and target RL environments, respectively. The predators did not share the rewards for simplicity.
@@
pre-trained 1M steps pre-trained 1M steps
DQN 0.04 $\pm$ 0.08 0.14 $\pm$ 0.08 4.09 $\pm$ 0.67 5.29 $\pm$ 0.86
DQfD 0.00 $\pm$ 0.00 0.16 $\pm$ 0.14 5.04 $\pm$ 0.39 5.08 $\pm$ 0.66
DQfAD 0.00 $\pm$ 0.00 0.30 $\pm$ 0.10 5.04 $\pm$ 0.39 5.10 $\pm$ 0.87
DQAS 0.42 $\pm$ 0.07 0.28 $\pm$ 0.08 5.00 $\pm$ 0.89 4.48 $\pm$ 0.81
DQAAS 0.42 $\pm$ 0.07 0.29 $\pm$ 0.12 5.00 $\pm$ 0.89 4.15 $\pm$ 0.87
@@
We then show the proportion of successful predation and DTW distance between the source and target trajectories for each model in Table 1. The results show that our approaches (DQAS and DQAAS) achieved better performances for both indicators than baselines. The obtained rewards and DTW distance had a trade-off relationship. In general, how to strike a balance is not obvious. In this task, with increased training steps, the DQAAS and DQAS first learned the ability to maximize a reward and then learned the reproducibility at the expense of the reward.
Here we show example results of the baseline (DQfD [hester2018deep]) and our approach (DQAAS) in Fig 3. The demonstration (left) shows that two predators chased the prey almost linearly and caught the prey in this scenario. In the source domain (demonstration), the predators were much faster than the prey (140 %), but in the target domain (RL), the predators were only slightly faster than the prey (107 %). Then, the task becomes more challenging than in the source domain and learning the Q-function correctly becomes more important to catch prey. Compared with the baseline, our approach correctly learned Q-function values in which the distribution concentrated near action 0 (here action 0 means moving toward the prey). These results imply the effectiveness of our approach quantitatively and qualitatively.
@@ 5.2 Performance on Football Tasks
Before using RL algorithm, we first created the demonstration dataset using real-world player location data in professional soccer league games. We used the data of 54 games in the Meiji J1 League 2019 season held in Japan. The dataset includes event data (i.e., labels of actions, e.g., passing and shooting, recorded at 30 Hz and the xy coordinates of the ball) and tracking data (i.e., xy coordinates of all players recorded at 25 Hz) provided by Data Stadium Inc. We extracted 218 last-pass-and-goal sequences and 1,385 last-pass sequences (including a ball lost) for training and pre-training of the RL model from the dataset. In pre-training, we split the dataset into 1,121 training, 125 validation, and 139 test sequences (or episodes). We set shot rewards ($+1$) for the attacker in addition to the above rewards and punishments (but we did not use the out-of-pitch punishment) because the goal reward was sparse and limited. In the target RL, we used 16 train and 5 test episodes from the above 218 episodes (we did not use the test condition in the target RL during the pre-training and training phases). We calculated and analyzed the obtained reward and DTW distance between the source and target trajectories in the test phase.
@@
pre-trained 0.5M steps pre-trained 5M steps
DQN 0.00 $\pm$ 0.00 0.52 $\pm$ 0.55 2.07 $\pm$ 0.48 1.71 $\pm$ 0.57
DQfD 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 3.27 $\pm$ 0.00 3.27 $\pm$ 0.00
DQfAD 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 3.27 $\pm$ 0.00 3.27 $\pm$ 0.00
DQAS 8.00 $\pm$ 0.00 8.00 $\pm$ 0.00 1.63 $\pm$ 0.00 1.63 $\pm$ 0.00
DQAAS 8.00 $\pm$ 0.00 8.00 $\pm$ 0.00 1.63 $\pm$ 0.00 1.63 $\pm$ 0.00
@@
DQN 0.00 $\pm$ 0.00 1.98 $\pm$ 2.43 2.35 $\pm$ 0.22 2.35 $\pm$ 0.26
CDS 0.08 $\pm$ 0.16 0.04 $\pm$ 0.07 2.37 $\pm$ 0.13 2.45 $\pm$ 0.56
DQfD 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 2.71 $\pm$ 0.00 3.68 $\pm$ 0.00
DQfAD 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 2.73 $\pm$ 0.00 2.67 $\pm$ 0.00
CDS+fD 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 2.88 $\pm$ 0.00 3.67 $\pm$ 0.00
CDS+fAD 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 2.88 $\pm$ 0.00 3.66 $\pm$ 0.00
DQAS 10.00 $\pm$ 0.00 10.00 $\pm$ 0.00 2.25 $\pm$ 0.00 2.25 $\pm$ 0.00
CDS+AS 10.00 $\pm$ 0.00 10.00 $\pm$ 0.00 2.24 $\pm$ 0.00 2.26 $\pm$ 0.00
DQAAS 10.00 $\pm$ 0.00 10.00 $\pm$ 0.00 2.25 $\pm$ 0.00 2.25 $\pm$ 0.00
CDS+AAS 10.00 $\pm$ 0.00 10.00 $\pm$ 0.00 2.24 $\pm$ 0.00 2.26 $\pm$ 0.00
@@ 6 CONCLUSION
We proposed a novel method for domain adaptation in RL from real-world multi-agent demonstration, which will bridge the gap between RL in cyberspace and data-driven modeling. In the experiments, using chase-and-escape and football tasks with the different dynamics between the unknown source and target environments, we showed that our approach balanced between the reproducibility and generalization more effectively than the baselines. In particular, we used the tracking data of professional football players as expert demonstrations in a football RL environment and demonstrated successful performances in both despite the larger gap between behaviors in the source and target environments than in the chase-and-escape task. Possible future research directions are to create a better multi-agent simulator and RL model utilizing domain knowledge for reproducing not only actions but also movements. In another direction, although modeling football movements would be currently challenging, for example, application to multi-animal behaviors will provide more scientifically valuable insights.

### 19 Dec 2023 v4

@@ 5.1 Performance on Chase-and-escape
First, we verified our method using a chase-and-escape task, in which the predators and prey interacted in a two-dimensional world with continuous space and discrete time. The numbers of predators and prey were 2 and 1, respectively. We first describe the common setting between the source (demonstration) and target RL tasks. The environment was constructed by modifying an environment called the predator-prey in Multi-Agent Particle Environment (MAPE) [lowe2017multi, tsutsui2022emergence, tsutsui2022collaborative]. Following [tsutsui2022collaborative], the play area size was constrained to the range of -1 to 1 on the $x$ and $y$ axes, all agent (predator/prey) disk diameters were set to 0.1, obstacles were eliminated, and predator-to-predator contact was ignored for simplicity. The predators were rewarded for capturing the prey ($+1$), namely contacting the disks, and punished for moving out of the area ($-1$), and the prey was punished for being captured by the predator or for moving out of the area ($-1$).
Fig. 2a shows an example of the chase-and-escape task. The time step was 0.1 s and the time limit in each episode was set to 30 s. The initial position of each episode was randomly selected from a range of -0.5 to 0.5 on the $x$ and $y$ axes. If the predator captured the prey within the time limit, the predator was successful; otherwise, the prey was successful. If one side (predators/prey) moved out of the area, the other side (prey/predators) was deemed successful. There are 13 actions including acceleration in 12 directions at every 30 degrees in relative coordinate system and doing nothing. For the relative mobility of predators and prey in both environments, to examine the effect of domain adaptation, we set the same mobility of the prey in the source and target RL environments, but for the predators, we set 120% and 110% of the prey mobility in the source and target RL environments, respectively. The predators did not share the rewards for simplicity.
@@
pre-trained 0.5M steps pre-trained 0.5M steps
DQN 0.04 $\pm$ 0.06 0.11 $\pm$ 0.03 4.12 $\pm$ 0.73 5.70 $\pm$ 0.62
DQfD 0.00 $\pm$ 0.00 0.04 $\pm$ 0.03 5.02 $\pm$ 0.34 4.94 $\pm$ 0.23
DQfAD 0.00 $\pm$ 0.00 0.06 $\pm$ 0.03 5.02 $\pm$ 0.34 4.80 $\pm$ 0.57
DQAS 0.25 $\pm$ 0.08 0.26 $\pm$ 0.08 5.37 $\pm$ 0.40 4.97 $\pm$ 1.16
DQAAS 0.25 $\pm$ 0.08 0.29 $\pm$ 0.09 5.37 $\pm$ 0.40 4.73 $\pm$ 1.07
@@
We then show the proportion of successful predation and DTW distance between the source and target trajectories for each model in Table 1. The results show that our approaches (DQAAS) achieved better performances for both indicators than baselines. The obtained rewards and DTW distance had a trade-off relationship. In general, how to strike a balance is not obvious. In this task, with increased training steps, the DQAAS first learned the ability to maximize a reward and then learned the reproducibility at the expense of the reward.
Here we show example results of the baseline (DQfD [hester2018deep]) and our approach (DQAAS) in Fig 3. The demonstration (left) shows that two predators chased the prey almost linearly and caught the prey in this scenario. In the source domain (demonstration), the predators were much faster than the prey (120 %), but in the target domain (RL), the predators were only slightly faster than the prey (110 %). Then, the task becomes more challenging than in the source domain and learning the Q-function correctly becomes more important to catch prey. Compared with the baseline, our approach correctly learned Q-function values in which the distribution concentrated near action 0 (here action 0 means moving toward the prey). These results imply the effectiveness of our approach quantitatively and qualitatively.
@@ 5.2 Performance on Football Tasks
Before using RL algorithm, we first created the demonstration dataset using real-world player location data in professional soccer league games. We used the data of 54 games in the Meiji J1 League 2019 season held in Japan. The dataset includes event data (i.e., labels of actions, e.g., passing and shooting, recorded at 30 Hz and the xy coordinates of the ball) and tracking data (i.e., xy coordinates of all players recorded at 25 Hz) provided by Data Stadium Inc. We extracted 198 last-pass-and-goal sequences and 1,385 last-pass sequences (including a ball lost) for training and pre-training of the RL model from the dataset. In pre-training, we split the dataset into 1,121 training, 125 validation, and 139 test sequences (or episodes). We set shot rewards ($+1$) for the attacker in addition to the above rewards and punishments (but we did not use the out-of-pitch punishment) because the goal reward was sparse and limited. In the target RL, we used 16 train and 5 test episodes from the above 198 episodes (we did not use the test condition in the target RL during the pre-training and training phases). We calculated and analyzed the obtained reward and DTW distance between the source and target trajectories in the test phase.
@@
pre-trained 0.5M steps pre-trained 0.5M steps
DQN 0.00 $\pm$ 0.00 1.40 $\pm$ 1.16 3.15 $\pm$ 0.47 2.69 $\pm$ 0.53
DQfD 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 4.58 $\pm$ 0.00 5.18 $\pm$ 0.00
DQfAD 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 4.58 $\pm$ 0.00 5.29 $\pm$ 0.01
DQAS 8.00 $\pm$ 0.00 8.00 $\pm$ 0.00 2.25 $\pm$ 0.00 2.25 $\pm$ 0.00
DQAAS 8.00 $\pm$ 0.00 8.00 $\pm$ 0.00 2.25 $\pm$ 0.00 2.25 $\pm$ 0.00
@@
DQN 0.00 $\pm$ 0.00 0.16 $\pm$ 0.13 3.22 $\pm$ 0.22 3.24 $\pm$ 0.22
CDS 0.12 $\pm$ 0.24 0.12 $\pm$ 0.24 3.10 $\pm$ 0.07 3.25 $\pm$ 0.12
DQfD 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 3.35 $\pm$ 0.00 3.78 $\pm$ 0.00
DQfAD 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 3.35 $\pm$ 0.00 4.28 $\pm$ 0.00
CDS+fD 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 3.71 $\pm$ 0.00 3.76 $\pm$ 0.00
CDS+fAD 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 3.71 $\pm$ 0.00 3.76 $\pm$ 0.00
DQAS 0.00 $\pm$ 0.00 6.00 $\pm$ 0.00 4.54 $\pm$ 0.00 3.27 $\pm$ 0.00
DQAAS 0.00 $\pm$ 0.00 6.00 $\pm$ 0.00 4.54 $\pm$ 0.00 3.30 $\pm$ 0.00
CDS+AS 6.00 $\pm$ 0.00 6.00 $\pm$ 0.00 3.25 $\pm$ 0.00 3.30 $\pm$ 0.00
CDS+AAS 6.00 $\pm$ 0.00 6.00 $\pm$ 0.00 3.25 $\pm$ 0.00 3.30 $\pm$ 0.00
@@ 6 CONCLUSION
We proposed a novel method for domain adaptation in RL from real-world multi-agent demonstration, which will bridge the gap between RL in cyberspace and data-driven modeling. In the experiments, using chase-and-escape and football tasks with the different dynamics between the unknown source and target environments, we showed that our approach balanced between the reproducibility and generalization more effectively than the baselines. In particular, we used the tracking data of professional football players as expert demonstrations in a football RL environment and demonstrated successful performances in both despite the larger gap between behaviors in the source and target environments than in the chase-and-escape task. Possible future research directions are to create a better multi-agent simulator and RL model utilizing domain knowledge for reproducing not only actions but also movements such as used by [tsutsui2023synergizing]. In another direction, although modeling football movements would be currently challenging, for example, application to multi-animal behaviors will provide more scientifically valuable insights.